Top to Bottom

New Results on Top and Bottom Quark Decays

Pekka K. Sinervo University of Toronto

1. Top and Bottom Quark Decays in the SM 2. Summary of Run I Data Samples

3. Top Quark Branching Fractions

4. Bottom Meson Hadronic Decays

5. Summary and Outlook for Run II HEP Seminar October 1996

Hadronic Bottom Decays

Start with

H = \frac{GF}{√2} Vcb \{[(d̄u) + (s̄c)] (c̄b)\}

• Renormalise H to take into account QCD corrections

Heff = \frac{GF}{√2} Vcb\{c1(μ) [(d̄u) + (s̄c)] (c̄b) +  (1)

c2(μ) [(c̄u)(d̄b) + (c̄c)(c̄b)] }

• For B → J/ψ decays, employ BSW factorisation:

dΓ(B+ → J/ψX) ∝ a2(mb)2|VcbV{∗ \atop cs}|2  (2)
|⟨J/ψ|(c̄c)|0⟩ |2 |⟨X|(s̄b)|B⟩|2

• Have to put in b quark form factors by hand

Model is fundamental enough to warrant detailed study

• CLEO/ARGUS have looked at many decays systematically

• But studies have had limitations – low statistics

– only produce B∘ & B+ (no B∘ s or baryons)

Hadron collider experiment overcomes some limitations

• Tevatron has produced ∼ 3 × 109 bb̄ events

Other Features of Top Decays

We have also looked for “exotic” decay modes

• t → Z∘q – Just run out of statistics

– Very weak limit on B(t → Z∘q)

Have also looked at polarisation of top decay

• May be just beyond reach statistically

Have yet to perform a complete study using all modes

• Incorporate hadronic modes

• Add in τ decay modes

Ratios of Branching Fractions

R{K∘ \atop K+} = \frac{fd}{fu} \frac{B(B∘ → J/ψK∘)}{B(B+ → J/ψK+ )} = 1.15 ± 0.27 ± 0.09,

R{K∗+ \atop K+} = \frac{B(B+ → J/ψK∗+ )}{B(B+ → J/ψK+ )} = 1.92 ± 0.60 ± 0.17,

R{K∗∘ \atop K+} = \frac{fd}{fu} \frac{B(B∘ → J/ψK∗∘)}{B(B+ → J/ψK+ )} = 1.59 ± 0.33 ± 0.12,

R{Kϕ \atop K+} = \frac{fs}{fu} \frac{B(B∘ s → J/ψϕ)}{B(B+ → J/ψK+ )} = 0.41 ± 0.12 ± 0.04,

R{K∗+ \atop K∘} = \frac{fu}{fd} \frac{B(B+ → J/ψK∗+ )}{B(B∘ → J/ψK∘)} = 1.68 ± 0.58 ± 0.11,
R{K∗∘ \atop K∘} = \frac{B(B∘ → J/ψK∗∘)}{B(B∘ → J/ψK∘)} = 1.39 ± 0.36 ± 0.10,
R{Kϕ \atop K∘} = \frac{fs}{fd} \frac{B(B∘ s → J/ψϕ)}{B(B∘ → J/ψK∘)} = 0.35 ± 0.12 ± 0.03,
R{K∗∘ \atop K∗+} = \frac{fd}{fu} \frac{B(B∘ → J/ψK∗∘)}{B(B+ → J/ψK∗+ )} = 0.83 ± 0.27 ± 0.07,

R{Kϕ \atop K∗+} = \frac{fs}{fu} \frac{B(B∘ s → J/ψϕ)}{B(B+ → J/ψK∗+ )} = 0.21 ± 0.08 ± 0.02,

R{Kϕ \atop K∗∘} = \frac{fs}{fd} \frac{B(B∘ s → J/ψϕ)}{B(B∘ → J/ψK∗∘)} = 0.26 ± 0.08 ± 0.02.